In a nornal distribution of pain estimates (0 = no pain , 100 = lots of pain) where m= 25 and s = 3 what percent would you expect to estimate their pain above a 32? please walk me through this problem
Answer #1

µ = 25

sd = 3

P(X > 32) = P(-4, 39-4

                         = P(Z > 32 – 25 3

                         = P(Z > 2.33)

                         = 1 - P(Z < 2.33)

                         = 1 - 0.9901

                         = 0.0099